a piece of equipment will function only when three components A,B and C are working. The probability of A failing during one year is 0.15, that of B is failing is 0.05 and that of C failing is 0.10. what is the probability that the equipment will fail before the end of the year?
1
Expert's answer
2020-10-27T20:07:14-0400

"P(failing)=1-P(not failing)"

=1-P(A not failing and B not failing and C not failing)

"=1-P(\bar{A}\bar{B}\bar{C})"

"=1-(0.85\times0.95\times0.9)=0.27325"